determine the mass of copper that will be deposited on a copper plate from an aqueous solution of copper(ii)tetraoxosulphate(vi) if the same current passed at the same time librated 10.8g of silver from an aqueous solution of silver trioxonitrate(v)[cu=64,Ag=108,Faraday's constant=96,500C mol]


1
Expert's answer
2021-01-25T04:33:20-0500

There are two ways to perform this task, one is the use of equivalent weights and the other is calculating the amount of current used in silver then using the current to calculate the mass of copper deposited.

The easiest is to use the former.


Now, according to Faraday's Second Law of electrolysis; when the same quantity of electricity is passed through several electrolytes, the mass of the substances deposited are proportional to their respective chemical equivalent or equivalent weights.

We know silver will undergo the reaction

Ag+ + e- "\to" Ag

Hence, its equivalent weight is "=\dfrac{RAM}{Valence} = \dfrac{108}{1} = 108"

Copper on the other hand will deposit as follow;

Cu2+ + 2e- "\to" Cu

and its equivalent weight "= \dfrac{64}{2} = 32"


From Faraday's second law;


"\dfrac{Eq.wt.of Ag}{Eq.wt.of Cu} = \dfrac{Mass. of Ag . deposited}{Mass .of.Cu.depoited}"


"\implies" "\dfrac{108}{32} = \dfrac{10.8g}{Mass. of.Cu. deposited}"


Mass of Cu deposited "= \dfrac{10.8g x 32}{108} = 3.2g" of copper was deposited